# Impact chronology of leftover planetesimals

### v1

@@ Abstract
Context. After the formation of the Moon the terrestrial planets were pummelled by impacts from planetesimals left over from terrestrial planet formation. Most lunar crater chronologies are fit with an exponentially decreasing impact rate with an e-folding time of about 150 Myr. Dynamical simulations consisting of leftover planetesimals and the asteroid belt should reproduce this impact rate.
@@
Methods. I ran dynamical simulations for 1 billion years using leftover planetesimals from the Grand Tack, Depleted Disc and Implantation models of terrestrial planet formation with the CPU version of the Gravitational ENcounters with GPU Acceleration (GENGA) N-body integrator. I fit the cumulative impacts on the Earth and Mars as well as the fraction of remaining planetesimals using a function that is a sum of exponentials with different weighing factors and e-folding times.
Results. Most fits require three or four terms. The fitted timescales cluster around million years (Myr), Myr, Myr and Myr. I attribute them to dynamical losses of planetesimals through different mechanisms: high-eccentricity Earth crossers and the secular resonance, Earth crossers, Mars crossers, and objects leaking on to Mars crossing orbits from beyond Mars. I place a constraint on the initial population using the known Archean terrestrial spherule beds, and I conclude that the Archean impacts were mostly created by leftover planetesimals. The inferred mass in leftover planetesimals at the time of the Moon’s formation was about 0.015 Earth masses.
Conclusions. The third time constant, , is comparable to that of modern crater chronologies. As such, the crater chronologies are indicative of impacts by an ancient population of Mars crossers. The initial perihelion distribution of the leftovers is a major factor in setting the rate of decline: to reproduce the current crater chronologies the number of Earth crossers at the time of the Moon’s formation had to be at most half of the Mars crossers. These results together place constraints on dynamical models of terrestrial planet formation.
@@
Planets and satellites: dynamical evolution and stability;1 Introduction
@@
In traditional dynamical models the terrestrial planets grew from a coagulation of planetesimals into protoplanets, which continued to remain submerged in a swarm of planetesimals for millions of years. This system eventually evolved into a giant impact phase, wherein the protoplanets collided with the planetesimals and with each other, which ultimately lead to the terrestrial planets. There exist many different dynamical simulations to explain this history (e.g. Chambers, 2001; Raymond et al., 2006, 2009; O’Brien et al., 2006; Walsh & Levison, 2019; Woo et al., 2021); a review of this can be found in Morbidelli et al. (2012), and a brief discussion of each model’s outcomes are given in Lammer et al. (2021).
@@
One key feature that all of the models share is that the terrestrial planets experience a protracted history of late accretion subsequent to planet formation. After core formation and initial separation of the silicate reservoirs (crust/mantle), left-over planetesimals on planet-crossing orbits were consumed by the terrestrial bodies as substantial mass supplements (Wetherill, 1977; Day et al., 2012). This took the form of impact bombardment by comets (Gomes et al., 2005), planetesimals remaining from primary accretion (Bottke et al., 2007), and asteroids from the main belt (Minton & Malhotra, 2010; Nesvorný et al., 2017) and the purported E-belt (Bottke et al., 2012). This late accretion thermally, structurally and chemically modified solid surfaces, most of which is directly visible as craters today.
@@
In the current Solar System, small objects that venture near the Earth – so called Near-Earth Objects, or NEOs – consist of asteroids and comets; the latter are mostly the Jupiter-family Comets (JFCs). The asteroids that venture near the Earth – the Near-Earth Asteroids, or NEAs – are typically injected onto Earth-crossing orbits from the main asteroid belt through the secular resonance near 2 au, the 3:1 Jovian mean-motion resonance near 2.5 au, and the Mars-crossing population (e.g. Gladman et al., 1997; Morbidelli & Gladman, 1998; Bottke et al., 2002; Granvik et al., 2018). Main belt asteroids are driven towards Mars-crossing orbits by numerous weak mean-motion, secular and three-body resonances (Migliorini et al., 1998; Nesvorný & Morbidelli, 1998; Morbidelli & Nesvorný, 1999), as well as the Yarkovsky effect (Farinella et al., 1998; Farinella & Vokrouhlický, 1999; Vokrouhlický & Farinella, 1998, 1999, 2000). From there they are rapidly driven to Earth-crossing orbits, and their population declines with a half life ranging from million years (Myr) for objects in the and 3:1 resonances (Gladman et al., 1997), to 60 Myr for the low-inclination Mars-crossers, to over 100 Myr for the high-inclination Mars-crossers (Migliorini et al., 1998; Michel et al., 2000; Bottke et al., 2002; Fernández & Helal, 2023).
@@
It is not obvious, however, that in the early Solar System these same timescales modulated the impact rate on the terrestrial planets. During this early epoch the impact rate was dominated by planetesimals left over from terrestrial planet formation, whose orbital distribution was very different from that of the main belt today (e.g Morbidelli et al., 2018; Brasser et al., 2020; Nesvorný et al., 2022). A constraint on the impact rate comes from crater chronology on the Moon and Mars, which for the first one billion years (Gyr) is fit well by a simple exponential decline with an e-folding time of 145-150 Myr for the Moon (Neukum et al., 2001; Werner et al., 2023), and 200 Myr for Mars (Werner et al., 2014; Werner, 2019). Even though the e-folding time of the Neukum and Werner chronologies are similar, the absolute crater density of the Werner chronology is much lower, resulting in the Werner chronology predicting ages up to 200 Myr older, and the latter requiring a much less massive impacting population. Dynamical simulations of leftover planetesimals, however, yield decline rates that vary greatly and are not necessarily consistent with that of the published crater chronologies (Morbidelli et al., 2018; Brasser et al., 2020; Nesvorný et al., 2022).
@@
The reason for these discrepancies could be that the initial conditions of the planetesimals at the time of the Moon-forming event were not consistent with the population that produced the current craters on the Moon and Mars. Both Morbidelli et al. (2018) and Brasser et al. (2020) relied on leftover planetesimals from terrestrial planet simulations in the framework of the Grand Tack model (Walsh et al., 2011), while Nesvorný et al. (2022) relied on results from the annulus model (Hansen, 2009). Yet none of these initial conditions appear very satisfactory in reproducing cratering rate advocated by the established crater chronologies and many different models of terrestrial planet formation are present in the literature. These models likely produce different leftover populations.
@@
In this work I investigate what the impact rate is of leftover planetesimals of three models of terrestrial planet formation simulations in my database, and whether these are or can be made consistent with modern crater chronologies. I also place constraints on what the orbital distribution of this population might have been, which of these models best reproduces this distribution, and what these constraints imply for models of terrestrial planet formation.
@@ 2 Numerical simulations: initial conditions and methods
This work relies on long-term simulations of massless test particles left over from simulations of terrestrial planet formation. Following Brasser et al. (2020), the initial conditions for this work were taken from a database of terrestrial planet formation simulations in my possession based on three models: the Grand Tack (GT) model (Walsh et al., 2011; Brasser et al., 2016), the Implantation (ABI) model (Brasser, 2025) and the Depleted Disc (DD) model (Izidoro et al., 2014; Mah & Brasser, 2021). I took a snapshot of the position and velocity vectors of the planetesimal population 60 Myr after the start of the simulations, which approximately coincides with 4.5 Ga. From these data I filtered out all planetesimals that had a perihelion distance au (2.1 au for DD), and an aphelion distance au so that they do not immediately venture near Jupiter. The initial values of semi-major axis (), eccentricity () and inclination () are shown in the top three panels of Fig. 1. The grey, orange, blue and red lines denote the required eccentricity to cross the orbits of Mercury, Venus, Earth and Mars, respectively. The colour coding is a proxy for the inclination, indicated by the colourbar. The bottom three panels show the cumulative distributions of perihelion, inclination and semi-major axis for all three models. All cumulative distributions are similar to either logistic or (offset) Rayleigh distributions. Specifically, the cumulative perihelion and inclination distributions are (offset) Rayleigh, and the semi-major axis distributions are logistic. These distributions are given by
@@
The GT model initially has the most Earth crossers (70% of all planetesimals), while the DD model has the least (only 30%). The ABI model has the highest average inclination, while this is lowest for the DD model. The planetesimals in the GT model are generally closer to the Sun than in the ABI and DD models, for which the cumulative semi-major axis distributions are very similar up to 2 au. Given these different initial conditions the expectation is that the dynamical evolution of all three populations will evolve somewhat differently.
The initial number of planetesimals from each set of terrestrial planet formation simulations was between 9 000 (9k) and 35 000 (35k), so I created clones of the initial test particle distribution using the KernelDensity Python module from the scikit-learn package (Pedregosa et al., 2011). A Gaussian kernel with a width of 0.005 was used on the semi-major axis, eccentricity and inclination; the other three angles were chosen uniformly random between 0 and 360∘. An example of this cloning in shown in Fig. 2. The black dots show the initial distribution, and the red dots are the clones. The width parameter was obtained through trial and error: I experimented with a larger width parameter, but I found that by increasing the width the eccentricity distribution of the clones, which was the most sensitive to the kernel width, became too wide.
@@
Simulations were run for 1 Gyr with a time step of 0.01 yr ( time steps in total). Each simulation began with 8192 test particles representing the leftover planetesimal population. Simulation data was written to disc every 100 000 years (100 kyr). Planetesimals were removed once they were further than 40 au from the Sun (whether bound or unbound), or when they collided with a planet or when they were closer than 0.05 au from the Sun. To keep simulation time reasonable I did not include corrections from General Relativity because this results in a 33% speed penalty. The code takes about 3-6 million steps per particle per second on Vega using four threads on AMD EPYC 7H12 CPUs clocked at 3.1 GHz; maximum speed was achieved when the number of test particles . I used two CPU threads when and only one when . For each terrestrial planet formation model I ran 16 simulations with a total of 131k test particles.
@@ 3.1 Global dynamical evolution
In Fig. 3 I show snapshots of the semi-major axis versus eccentricity for the ABI simulations, where the colour coding of the planetesimals is a proxy for their inclination, and the panel titles refer to the simulation time in million years (Myr). After 400 Myr almost all Earth and Venus crossers are eliminated, but there is still a substantial population of Mars crossers. There is also a small group of planetesimals that are seemingly stable between the Earth and Mars for 1 Gyr. This is an artefact of the simulations because these bodies would have had their orbits modified by the Yarkovsky effect and therefore this region is expected to be empty – as it is today. Simulations of these same initial conditions including the Yarkovsky effect are ongoing and will be discussed in future works. Last, there is a long-term stable population at moderate inclination between Mars and the resonance, which is located at 2.05 au at low inclination (Williams & Faulkner, 1981); this population is akin to the Hungaria family (Bottke et al., 2012). Morbidelli & Nesvorný (1999) showed that bodies beyond the resonance may slowly leak onto Mars-crossing orbits due to chaotic motion induced by resonance overlap. In addition, some bodies remained on stable orbits in the main belt between 2.1 and 3 au.
@@
A different visualisation of the evolution of the planetesimals is shown in Fig. 4, which is a heat map of the semi-major axis and perihelion distance of all bodies in all simulations for each model. The top panel is for GT, the middle for ABI, and the bottom for DD. The gradient increases from blue to red, indicated by the colourbar, with blue dots indicating few planetesimals ventured in this region of phase space, and red indicating that many remained in another region. The terrestrial planets plus Vesta and Ceres are shown with large bullets. These heat maps just show the counts per bin, and not the residence time though it is easy to convert between the two. It is immediately clear that the maps for the ABI and DD models show some similarities, while the map for the GT model is rather different.
@@
The distributions shown in the heat maps depend on the initial conditions for the models. In the ABI and DD models, there is a sharp increase in density, or residence time, for planetesimals that are Earth-crossing versus those that are not Earth-crossing, as indicated by the two red triangular regions in these panels. In the ABI model, the first region appears at au and au, while the second one is in the DD model and has au and au. The boundaries in semi-major axis for the second region are marked by the resonance at 2 au and the 3:1 Jovian resonance at 2.5 au. In both plots the E-belt region between 1.7 au and 2 au, as proposed by Bottke et al. (2012), remains occupied for a long time. In the DD model the inner main belt is heavily populated, a consequence of the initial conditions. In contrast, the GT heat map looks substantially different, with the reddest region having au and au apart from a small sample with au and . The GT model has the greatest proportion of initial Earth-crossers, so that it is expected that this population has a shorter average lifetime than those of the ABI and DD models. The heat maps suggest that the initial perihelion distribution plays an important role in the rate of decline of the population, and thus in the impact rate onto the inner planets.
@@
In Fig. 5 I present a snapshot of the semimajor axis-inclination distribution at the end of the simulations for all three different cases, and compare them to the actual main belt in the bottom right panel. It is clear that the DD case fares poorly in that it only produces the Hungaria group near au and and the inner main belt underneath the secular resonance for which au, but not the high inclination population above the . The GT and ABI cases populate all of the main groups to some extent, except the one between 3 au and 3.25 au. The ABI model leaves more mass in the Hungaria group than the GT case, while GT populates the middle main belt between 2.5 au and 2.8 au better than ABI does. Overall the GT and ABI models fare better than DD.

### v3

@@ Abstract
Context. After the formation of the Moon, the terrestrial planets were pummelled by impacts from planetesimals left over from terrestrial planet formation. Most lunar crater chronologies are fitted with an exponentially decreasing impact rate with an e-folding time of about 150 Myr. Dynamical simulations consisting of leftover planetesimals and the asteroid belt should reproduce this impact rate.
@@
Methods. I ran dynamical simulations for 1 billion years using leftover planetesimals from the grand tack, depleted disc, and implantation models of terrestrial planet formation with the CPU version of the Gravitational ENcounters with GPU Acceleration (GENGA) N-body integrator. I fitted the cumulative impacts on the Earth and Mars as well as the fraction of remaining planetesimals using a function that is a sum of exponentials with different weighing factors and e-folding times.
Results. Most fits require three or four terms. The fitted timescales cluster around million years (Myr), Myr, Myr, and Myr. I attribute them to dynamical losses of planetesimals through different mechanisms: high-eccentricity Earth-crossers and the secular resonance, Earth-crossers, Mars-crossers, and objects leaking onto Mars-crossing orbits from beyond Mars. I placed a constraint on the initial population using the known Archean terrestrial spherule beds, and I conclude that the Archean impacts were mostly created by leftover planetesimals. The inferred mass in leftover planetesimals at the time of the Moon’s formation was about 0.015 Earth masses.
Conclusions. The third time constant, , is comparable to that of modern crater chronologies. As such, the crater chronologies are indicative of impacts by an ancient population of Mars-crossers. The initial perihelion distribution of the leftovers is a major factor in setting the rate of decline: to reproduce the current crater chronologies, the number of Earth-crossers at the time of the Moon’s formation had to be at most half that of the Mars-crossers. These results together place constraints on dynamical models of terrestrial planet formation.
@@
Planets and satellites: dynamical evolution and stability; Planets and satellites: terrestrial planets; Minor planets, asteroids: general1 Introduction
@@
In traditional dynamical models, the terrestrial planets grew from a coagulation of planetesimals into protoplanets, which remained submerged in a swarm of planetesimals for millions of years. This system eventually evolved into a giant impact phase, wherein the protoplanets collided with the planetesimals and with each other, which ultimately led to the terrestrial planets. There exist many different dynamical simulations to explain this history (e.g. Chambers, 2001; Raymond et al., 2006, 2009; O’Brien et al., 2006; Walsh & Levison, 2019; Woo et al., 2021); a review can be found in Morbidelli et al. (2012), and a brief discussion of each model’s outcomes is given in Lammer et al. (2021).
@@
One key feature that all of the models share is that the terrestrial planets experienced a protracted history of late accretion subsequent to planet formation. After core formation and the initial separation of the silicate reservoirs (e.g. the crust and mantle), leftover planetesimals on planet-crossing orbits were consumed by the terrestrial bodies as mass supplements (Wetherill, 1977; Day et al., 2012). This took the form of impact bombardment by comets (Gomes et al., 2005), planetesimals remaining after the primary phase of accretion (Bottke et al., 2007), and asteroids from the main belt (Minton & Malhotra, 2010; Nesvorný et al., 2017) and the purported E-belt (Bottke et al., 2012). This late accretion thermally, structurally, and chemically modified solid surfaces, most of which is directly visible as craters today.
@@
In the current Solar System, small objects that venture near the Earth – so called near-Earth objects – consist of asteroids and comets; the latter are mostly the Jupiter-family comets. The asteroids that venture near the Earth – the near-Earth asteroids – are typically injected onto Earth-crossing orbits from the main asteroid belt through the secular resonance near 2 au, the 3:1 Jovian mean-motion resonance near 2.5 au, and the Mars-crossing population (e.g. Gladman et al., 1997; Morbidelli & Gladman, 1998; Bottke et al., 2002; Granvik et al., 2018). Main belt asteroids are driven towards Mars-crossing orbits by numerous weak mean-motion, secular, and three-body resonances (Migliorini et al., 1998; Nesvorný & Morbidelli, 1998; Morbidelli & Nesvorný, 1999), as well as the Yarkovsky effect (Farinella et al., 1998; Farinella & Vokrouhlický, 1999; Vokrouhlický & Farinella, 1998, 1999, 2000). From there they are rapidly driven to Earth-crossing orbits, and their population declines with a half life ranging from Myr for objects in the and 3:1 resonances (Gladman et al., 1997) to 60 Myr for the low-inclination Mars-crossers, to over 100 Myr for the high-inclination Mars-crossers (Migliorini et al., 1998; Michel et al., 2000; Bottke et al., 2002; Fernández & Helal, 2023).
@@
It is not obvious, however, whether in the early Solar System these same timescales modulated the impact rate on the terrestrial planets. During this early epoch the impact rate was dominated by planetesimals left over from terrestrial planet formation, whose orbital distribution was very different from that of the main belt today (e.g. Morbidelli et al., 2018; Brasser et al., 2020; Nesvorný et al., 2022). A constraint on the impact rate comes from crater chronology on the Moon and Mars, which for the first 1 Gyr is fitted well by a simple exponential decline with an e-folding time of 145-150 Myr for the Moon (Neukum et al., 2001; Werner et al., 2023) and 200 Myr for Mars (Werner et al., 2014; Werner, 2019). Even though the e-folding time of the Neukum and Werner chronologies are similar, the absolute crater density of the Werner chronology is much lower, resulting in the Werner chronology predicting ages of up to 200 Myr older; this slower decline requires a much less massive impacting population. Dynamical simulations of leftover planetesimals, however, yield decline rates that vary greatly and are not necessarily consistent with that of the published crater chronologies (Morbidelli et al., 2018; Brasser et al., 2020; Nesvorný et al., 2022).
@@
The reason for these discrepancies could be that the initial conditions of the planetesimals at the time of the Moon-forming event were not consistent with the population that produced the current craters on the Moon and Mars. Both Morbidelli et al. (2018) and Brasser et al. (2020) relied on leftover planetesimals from terrestrial planet simulations in the framework of the grand tack (GT) model (Walsh et al., 2011), while Nesvorný et al. (2022) relied on results from the annulus model (Hansen, 2009). Yet none of these initial conditions appear to reproduce the cratering rate advocated by the established crater chronologies with a good degree of satisfaction, and many different models of terrestrial planet formation are present in the literature. These models likely produce different leftover populations.
@@
In this work I determine the impact rate of leftover planetesimals of three models of terrestrial planet formation simulations in my database, and whether these are or can be made consistent with modern crater chronologies. I also place constraints on what the orbital distribution of this population might have been, which of these models best reproduces this distribution, and what these constraints imply for models of terrestrial planet formation.
@@ 2 Numerical simulations: initial conditions and methods
This work relies on long-term simulations of massless test particles left over from simulations of terrestrial planet formation. Following Brasser et al. (2020), the initial conditions for this work were taken from a database of terrestrial planet formation simulations in my possession based on three models: the GT model (Walsh et al., 2011; Brasser et al., 2016), the (asteroid belt) implantation (ABI) model (Brasser, 2025), and the depleted disc (DD) model (Izidoro et al., 2014; Mah & Brasser, 2021). Briefly, the GT model relies on the inward-outward migration of Jupiter and Saturn to truncate the planetesimal disc near 1 au to keep Mars’ mass low. The DD model artificially lowers the surface density of solids beyond 1-1.5 au to starve the region near Mars of material to accrete. The ABI model relies on the implantation of planetesimals near the gas giants to account for terrestrial accretion of carbonaceous chondrite-like material (e.g. Dauphas, 2017). I took a snapshot of the position and velocity vectors of the planetesimal population 60 Myr after the start of the simulations, which approximately coincides with 4.5 Ga. From these data I filtered out all planetesimals that had a perihelion distance au (2.1 au for DD to establish how the inner main belt affects the decline), and an aphelion distance au so that they do not immediately venture near Jupiter. The initial values of semi-major axis (), eccentricity () and inclination () are shown in the top three panels of Fig. 1. The bottom three panels show the cumulative distributions of perihelion, inclination, and semi-major axis for all three models. All cumulative distributions are similar to either logistic or (offset) Rayleigh distributions. Specifically, the cumulative perihelion and inclination distributions are (offset) Rayleigh, and the semi-major axis distributions are logistic. These distributions are given by
@@
The GT model initially has the most Earth-crossers (70% of all planetesimals), while the DD model has the least (only 30%). The ABI model has the highest average inclination, while this is lowest for the DD model. The planetesimals in the GT model are generally closer to the Sun than in the ABI and DD models, for which the cumulative semi-major axis distributions are very similar up to 2 au. Given these different initial conditions the expectation is that the dynamical evolution of all three populations will evolve somewhat differently.
The initial number of planetesimals from each set of terrestrial planet formation simulations was between 9 000 (9k) and 35 000 (35k), so I created clones of the initial test particle distribution using the KernelDensity Python module from the scikit-learn package (Pedregosa et al., 2011). A Gaussian kernel with a width of 0.005 was used on the semi-major axis, eccentricity and inclination; the other three angles were chosen uniformly random between 0 and 360∘. An example of this cloning in shown in Fig. 2. The width parameter was obtained through trial and error: I experimented with a larger width parameter, but I found that by increasing the width the eccentricity distribution of the clones, which was the most sensitive to the kernel width, became too wide.
@@
Simulations were run for 1 Gyr with a time step of 0.01 yr ( time steps in total). Each simulation began with 8192 test particles representing the leftover planetesimal population. Simulation data were written to disc every 100 000 years (100 kyr). Planetesimals were removed once they were further than 40 au from the Sun (whether bound or unbound), or when they collided with a planet or when they were closer than 0.05 au from the Sun. To keep simulation time reasonable I did not include corrections from General Relativity because this results in a 33% speed penalty. The code takes about 3-6 million steps per particle per second on Vega using four threads on AMD EPYC 7H12 CPUs clocked at 3.1 GHz; maximum speed was achieved when the number of test particles . I used two CPU threads when and only one when . For each terrestrial planet formation model I ran 16 simulations with a total of 131k test particles.
@@ 3.1 Global dynamical evolution
In Fig. 3 I show snapshots of the semi-major axis versus eccentricity for the ABI simulations. After 400 Myr almost all Earth and Venus crossers are eliminated, but there is still a substantial population of Mars-crossers. There is also a small group of planetesimals that are seemingly stable between the Earth and Mars for 1 Gyr. This is an artefact of the simulations because these bodies would have had their orbits modified by the Yarkovsky effect and therefore this region is expected to be empty – as it is today. Simulations of these same initial conditions including the Yarkovsky effect are ongoing and will be discussed in future works. Last, there is a long-term stable population at moderate inclination between Mars and the resonance, which is located at 2.05 au at low inclination (Williams & Faulkner, 1981); this population is akin to the Hungaria family (Bottke et al., 2012). Morbidelli & Nesvorný (1999) showed that bodies beyond the resonance can slowly leak onto Mars-crossing orbits due to chaotic motion induced by resonance overlap. In addition, some bodies began and remained on long-term meta-stable orbits in the main belt with semi-major axis between 2.1 and 3 au and perihelion distance au.
@@
A different visualisation of the evolution of the planetesimals is shown in Fig. 4, which is a heat map of the semi-major axis and perihelion distance of all bodies in all simulations for each model. The terrestrial planets plus Vesta and Ceres are shown with large bullets. These heat maps show the counts per bin and not the residence time, though it is easy to convert between the two. It is immediately clear that the maps for the ABI and DD models show some similarities, while the map for the GT model is rather different.
@@
The distributions shown in the heat maps depend on the initial conditions for the models. In the ABI and DD models, there is a sharp decrease in density, or residence time, for planetesimals that are Earth-crossing versus those that are not Earth-crossing, as indicated by the two red triangular regions in these panels. In both the ABI and DD models there is one such region with au and au, while in the DD model the second one has au and au. The second region is in the DD model. It is more heavily populated than in the same regions in the ABI and GT models due to the higher initial perihelion cutoff and the higher implantation for the DD model. The boundaries in semi-major axis for the second region are marked by the resonance at 2 au and the 3:1 Jovian resonance at 2.5 au. In both plots the E-belt region between 1.7 au and 2 au, as proposed by Bottke et al. (2012), remains occupied for a long time. In the DD model the inner main belt is heavily populated, a consequence of the initial conditions. In contrast, the GT heat map looks substantially different, with the reddest region having au and au apart from a small sample with au and . The GT model has the greatest proportion of initial Earth-crossers, so that it is expected that this population has a shorter average lifetime than those of the ABI and DD models. The GT heat map also has the orange region extend all the way down to Mercury, while in the ABI and DD models it stops at Venus. The reason for this difference is that the initial perihelion distribution for the GT model extends all the way down to 0.1 au while this is 0.3 au for the ABI and DD models. One immediate outcome of this is that the impact probability with Mercury in the GT model is a factor of three higher than for the other two models. The heat maps suggest that the initial perihelion distribution plays an important role in the rate of decline of the population, and thus in the impact rate onto the inner planets.
@@
In Fig. 5 I present a snapshot of the semi-major axis-inclination distribution at the end of the simulations for all three different cases, and compare them to the actual main belt in the bottom right panel. Even though the initial conditions used here all came from terrestrial planet formation models, all of these models leave some material in the main belt region. Comparing its orbital distribution with that of the present main belt serves as a constraint of the validity of those models, and associated predictions for the long-term bombardment of the terrestrial planets from the main belt. It is clear that the DD case fares poorly in that it only produces the Hungaria group near au and and the inner main belt underneath the secular resonance for which au, but not the high inclination population above the . The GT and ABI cases populate all of the main groups to some extent, except the one between 3 au and 3.25 au. The ABI model leaves more mass in the Hungaria group than the GT case, while GT populates the middle main belt between 2.5 au and 2.8 au better than ABI does. Overall the GT and ABI models fare better than DD.